# Adaptive Hyper-Graph Convolution Network for Skeleton-based Human Action Recognition with Virtual Connections

### v2

@@ Abstract
The shared topology of human skeletons motivated the recent investigation of graph convolutional network (GCN) solutions for action recognition. However, most of the existing GCNs rely on the binary connection of two neighboring vertices (joints) formed by an edge (bone), overlooking the potential of constructing multi-vertex convolution structures. Although some studies have attempted to utilize hyper-graphs to represent the topology, they rely on a fixed construction strategy, which limits their adaptivity in uncovering the intricate latent relationships within the action. In this paper, we address this oversight and explore the merits of an adaptive hyper-graph convolutional network (Hyper-GCN) to achieve the aggregation of rich semantic information conveyed by skeleton vertices. In particular, our Hyper-GCN adaptively optimises the hyper-graphs during training, revealing the action-driven multi-vertex relations. Besides, virtual connections are often designed to support efficient feature aggregation, implicitly extending the spectrum of dependencies within the skeleton. By injecting virtual connections into hyper-graphs, the semantic clues of diverse action categories can be highlighted. The results of experiments conducted on the NTU-60, NTU-120, and NW-UCLA datasets demonstrate the merits of our Hyper-GCN, compared to the state-of-the-art methods. Specifically, we outperform the existing solutions on NTU-120, achieving 90.5% and 91.7% in terms of the top-1 recognition accuracy on X-Sub and X-Set.
Index Terms:
Action recognition, Video Understanding, Computer Vision, Graph Learning.
I Introduction
Skeleton-based human action recognition is a popular research topic in artificial intelligence, with practical applications in video understanding, video surveillance, human-computer interaction, robot vision, VR, and AR [1, 2, 3, 4, 5, 6, 7]. In general, a skeleton sequence contains a series of 2D or 3D coordinates, which can easily be collected by low-cost depth sensors or obtained by video-based pose estimation algorithms. Compared to RGB and optical flow images, skeleton data, which represents the basic physical structure of a human being, is of lower dimension, conveying human action with higher efficiency. Moreover, it is robust to illumination changes and scene variations. For these reasons, the adoption of this structural data is very popular in skeleton-based action recognition [8, 9, 10, 11, 12, 13, 14].
@@ I Introduction
Figure 2: The framework of our Hyper-GCN. Part (a) represents the Multi-Scale Hyper-GC (MS-HGC) module, where channels are divided into 8 branches. Part (b) represents the process of constructing an adaptive hyper-graph. Part (c) represents the architecture of Hyper-GCN. $\mathbf{F}_{h}$ represents the hyper-joints which are learnable parameters. $\hat{A}_{p}$ represents the physical topology. $HA$ represents The fused topology of the adaptive hyper-graph and the human physical structure.
II Related Work
II-A Graph Topology for Action Recognition
@@
II-B Feature Configurations for Action Recognition
@@ III-A Preliminaries
III-B Hyper-graph
Here, we use $\mathcal{G_{H}}=(\mathcal{V_{H}},\mathcal{E_{H}},\mathcal{W_{H}})$ to define the spatial hyper-graph [31] with human skeleton. $\mathcal{V_{H}}$ and $\mathcal{E_{H}}$ follow similar definitions in the normal graph, which represent the set of joints and the set of hyper-edges. In addition, we introduce $\mathcal{W_{H}}$ to represent the weights of each hyper-edge. Since a hyper-edge contains multiple joints, the corresponding topology can no longer be simply represented by an adjacency matrix. Therefore, we introduce the incidence matrix $H\in\mathbb{R}^{N\times M}$ to describe the topology of each joint in the hyper-graph. $N$ represents the number of joints and $M$ represents the number of hyper-edges. Given $v\in\mathcal{V_{H}}$ and $e\in\mathcal{E_{H}}$, the values of incidence matrix can be determined by:
@@ III-B Hyper-graph
III-C Adaptive Non-uniform Hyper-graph
@@
Given the general feature $\mathbf{X}\in\mathbb{R}^{N\times C}$ of spatial skeleton joints, $N$ represent the number of joints, $C$ represent the channels. Our A-NHG construct the hyper-edge with each joint as the center of mass, and a total of $N$ hyper-edges are constructed. A-NHG utilises the Euclidean distance to measure the joint difference.
@@
$h_{i,j}$ =$=\begin{cases}\frac{\text{exp}(-m_{i,j})}{\sum_{k\in set_{i}}{\text{exp}(-m_{i,k})}},&j\in set_{i},\\ 0,&j\notin set_{i}\end{cases}$ (8)
@@
III-D Multi-head Hyper-graph Convolution
@@
We use separate branches to independently perform hyper-graph convolution on the topologies represented by the multi-head hyper-graphs, thereby enhancing the computational efficiency of the Hyper-GCN. For instance, the feature $\mathbf{F_{in}\in\mathbb{R}^{C_{in}\times T\times V}}$ is divided into $8$ separate branches along the channel dimension, which are processed by $8$ M-HGC parallelly, delivering $8$ separate hyper-graphs.
Next, we perform temporal pooling, and then the obtained spatial information is the evidence to determine the optimal hyper-graph construction. This can efficiently decouple the temporal and spatial clues. After performing temporal average-pooling, we can obtain $\mathbf{\bar{F}_{in}}\in\mathbb{R}^{C_{in}\times V}$.
The hyper-graph is depicted by the incidence matrix and weight matrix. So, we use a separate mapping function for each head to embed the features into the subspace for constructing the hyper-graph. This design enables preserving the original features to reflect their spatial relevance. After that, we introduce the A-NHG to obtain the incidence matrix. For the weight matrix, we adopt the Channel Attention [54]. The hyper-graph obtained operation can be formulated as follows:
$H$ $=\uplus^{8}_{i=1}\text{A-NHG}(\mathbf{\bar{F}^{k}_{in}}),$ (9)
$W$ $=\sigma^{2}(\Psi^{2}(\uplus^{8}_{i=1}\sigma^{1}(\Psi^{1}_{i}(\mathbf{\bar{F}^{k}_{in}})))),$ (10)
@@
$\mathbf{F_{out}}$ $=\uplus^{8}_{i=1}{(\hat{A_{i}}+\alpha\cdot\hat{H_{i}})\mathbf{F^{k}_{in}}P_{i}},$ (11)
$\hat{A}$ and $\hat{H}$ denote the normalized physical adjacency matrix and normalized incidence matrix. $P$ is the learnable weight parameter for feature transform. $\alpha$ is a learnable parameter for the topology fusion of each head.
Category Methods Publication Modalities NTU-RGB+D 60 NTU-RGB+D 120 NW-UCLA (%)
X-Sub (%) X-View (%) X-Sub (%) X-Set (%)
GCN-based MST-GCN [40] AAAI21 J+B+JM+BM 91.5 96.6 87.5 88.8 -
CTR-GCN [13] ICCV21 J+B+JM+BM 92.4 96.4 88.9 90.6 96.5
EfficientGCN-B4 [49] TPAMI22 J+B+JM+BM 91.7 95.7 88.3 89.1 -
InfoGCN [12] CVPR22 J+B+JM+BM 92.7 96.9 89.4 90.7 96.6
FR Head [55] CVPR23 J+B+JM+BM 92.8 96.8 89.5 90.9 96.8
HD-GCN* [19] ICCV23 J+B+J’+B’ 93.0 97.0 89.8 91.2 96.9
DS-GCN [14] AAAI24 J+B+JM+BM 93.1 97.5 89.2 90.3 -
BlockGCN [20] CVPR24 J+B+JM+BM 93.1 97.0 90.3 91.5 96.9
HGCN-based Hyper-GNN [32] TIP21 J+B+JM+BM 89.5 95.7 - - -
Selective-HCN [33] ICMR22 J+B+JM+BM 90.8 96.6 - - -
DST-HCN [56] ICME23 J+B+JM+BM 92.3 96.8 88.8 90.7 96.6
Hyper-GCN (w/o ensemble) - J 91.4 95.5 87.0 88.7 95.5
Hyper-GCN (4-ensemble) - J+B+JM+BM 93.4 97.1 90.5 91.7 97.1
TABLE I: Comparison of Hyper-GCN with advanced solutions on NTU-RGB+D 60, NTU-RGB+D 120, and NW-UCLA datasets. For a fair comparison, we use a 4-streams ensemble to evaluate the performance of the mentioned methods. J, B, JB, and JM represent the joint, bone, joint motion, and bone motion. In addition, for HD-GCN [19], we use the results from 4-streams according to the original papers for a fair comparison. The bold font and the underline font represents the best result and 2-nd best result.
III-E Virtual Connections
@@
The shape of hyper-joints is consistent with the feature $\mathbf{F_{p}}\in\mathbb{R}^{C\times T\times V}$, which is described as $\mathbf{F_{h}}\in\mathbb{R}^{C\times T\times V_{h}}$. Specifically, the learning of hyper-joints is supervised by the loss function, which guides these hyper-joints to support generalisable driven features embedded within large amounts of data. Furthermore, we set independent hyper-joints at each layer of Hyper-GCN, aiming to harmonise features at different depths. Typically, these hyper-joints are involved in spatial hyper-graph convolution rather than in temporal convolution.
@@
$C=\frac{\mathbf{F_{h}F^{T}_{h}}}{||\mathbf{F_{h}}||^{2}},$ (12)
Specifically, in Divergence Loss, we measure the differences of hyper-joints with the mean of the cosine matrix $C$ in each layer. Since the correlation of hyper-joints themselves cannot be optimised, we manually eliminate this part by minus the $V_{h}$. The calculation can be formulated as:
@@
where $V_{h}$ represents the number of hyper-joints we introduced. $C_{l}$ represents the cosine matrix of the $l$-th layer.
Additionally, we manually connect the hyper joints to all the physical joints, which optimises the topology from the datasets.
III-F Entire Architecture
Hyper-GCN consists of an embedding layer [12], 9 spatial-temporal convolution layers as shown in Figure 2 (c). Each layer consists of the proposed M-HGC and Multi-Scale Temporal convolution (MS-TC) [13]. 9 Layers are categorised into 3 stages with Dense connections. The channels in each stage are set to 128, 256 and 512.
IV Evaluation
IV-A Datasets
@@
IV-B Implementation Details
Our training and evaluation stages are on a single RTX 3090 GPU. In the training phase, Hyper-GCN is optimised by Stochastic Gradient Descent (SGD) with Nesterov momentum set at 0.9 and a weight decay at 0.0004. Our implementation uses label smooth cross-entropy loss with the Divergence Loss we proposed. We set a total of 140 epochs with the start 5 warm-up epochs. The initial learning rate is 0.05, which is reduced to 0.005 at epoch 110 and to 0.0005 at epoch 120.
IV-C Comparison with the State-of-the-Art
Multi-stream ensembles proposed in  [57] have been proven effective by most of the existing state-of-the-art methods. We also use a 4-streams ensemble to evaluate our model. The detailed results are reported in Table I. To the best of our knowledge, a comparison with SOTA shows that our proposed Hyper-GCN achieves the SOTA level on each dataset. This demonstrates the effectiveness and merits of Hyper-GCN.
IV-D Ablation Study
We conduct ablation experiments with visualisations to analyse the effectiveness of our design. All the ablation study is evaluated with the joint modality.
@@ IV-D Ablation Study
TABLE II: Ablation of the hyper-parameter $K$ in A-NHG. The Uniform represents K-uniform hyper-graph, in which one hyper-edge contains K joints. The non-uniform represents non-uniform hyper-graph, in which one joint is contained by $K$ hyper-edge.
Figure 4: Visualisation of hyper-graph in selected actions. The size of a joint means its weight within the hyper-edge. For ease of observation, we visualise the hyper-edges with the larger weight constructed in each sample.
@@
The results are reported in Table II. In both uniform and non-uniform hyper-graphs, a very large $K$ means that more joints are contained in each hyper-edge. Even though these ”extra” joints may be optimised for minimal weight, they are likely to act as ”noise”, disrupting the information interaction represented by the hyper-edge. As the $K$ decreases, it is difficult for the hyper-edge to represent complex joint combination patterns. Therefore, setting $K=5$ in uniform and $K=9$ in Non-uniform achieves the best performance. The non-uniform hyper-graph of A-NHG outperforms the uniform hyper-graph, as the combination of skeleton joints in different actions is diverse.
Furthermore, we visualise the hyper-graph inferred by Hyper-GCN on selected actions in Figure 4. The hyper-graph construction focuses on the motion joints that are most relevant to the corresponding action category. For “Kicking something” as part (d), the right leg joints and the left hand joint are modeled as a hyper-edge. Because in this action, the person’s right hand previously extends, and the right leg swings relatively to the right hand. Similarly, in the action ”Stand up” as shown in part (b), most joints of the upper body move relatively to right foot joint, which is not moving in the first half of frames. In addition, ”Taking a selfie” and ”Drinking water” are also consistent with the above view as Figure 4 (a), (c).
@@
w/o M-HGC w/ 1 84.9 ($\uparrow 0.2$) 84.9 ($\uparrow 0.2$)
w/ 3 84.9 ($\uparrow 0.2$) 85.2 ($\uparrow 0.5$)
w/ 5 84.7 85.0 ($\uparrow 0.3$)
w M-HGC w/ 1 86.7 ($\uparrow 2.0$) 86.7 ($\uparrow 2.0$)
w/ 3 86.6 ($\uparrow 1.8$) 86.9 ($\uparrow 2.2$)
w/ 5 86.6 ($\uparrow 1.8$) 84.8 ($\uparrow 2.1$)
TABLE III: The ablations on hyper joints. w/ $N$ represents the number of hyper joints in each layer. w/ $\mathcal{L}_{h}$ represents the training with Divergence Loss. The hyper-parameter $K$ of M-HGC is set to $9$.
@@
Virtual Connections: We conduct the ablation study to analyse the hyper joints and the Divergence Loss. The results are listed in Table III. By comparison, the performance is not monotonically increasing with the number of hyper joints. The case of introducing only $3$ hyper joints achieves the best performance. As our hyper joints are learned from a large amount of data, involving a large number of hyper joints can introduce redundant and ambiguous clues, degrading the performance.
@@
Methods Publications NTU RGB+D 120 GFLOPs Params
X-Sub(%) X-Set(%)
CTR-GCN[13] ICCV21 88.9 90.4 1.97 1.46
InfoGCN[12] CVPR23 89.4 90.4 1.70 1.52
FR-Head[55] CVPR23 89.5 90.9 2.0 -
BlockGCN[20] CVPR24 90.3 91.5 1.63 1.31
Hyper-GCN (-) - 90.4 91.5 1.63 1.06
DSTA-Net[25] ACCV20 86.6 89.0 16.18 3.45
IIP-Transformer[26] BigData23 89.9 90.9 7.20 2.90
SkateFormer[44] ECCV24 89.8 91.4 3.62 2.03
Hyper-GCN - 90.5 91.7 2.88 2.34
TABLE IV: The comparison of the performance and complexity. Hyper-GCN (-) represents the lightweight version.
IV-E Comparison of Complexity
Increasing the number of channels provides considerable improvement, which also increases the complexity to some extent. Given that advanced lightweight GCN-based methods are between 1M and 2M. For a fair comparison, we set channels in the last 3 layers to 256 to reduce the parameters to 1.06M, which is lighter weight than all involved counterparts. The performance of our lightweight version compared with GCN-based methods is reported in Table IV. In fact, on NTU RGB+D 120, which is the most challenge dataset, the performance of Hyper-GCN is degraded, but still achieves the SOTA level with the minimal parameter volume. In addition, on a comparable basis with the parameters of Transformer-based methods, our Hyper-GCN also outperform the SOTA.
V Conclusion
In this paper, we propose an Adaptive Hyper-graph Convolutional Network for skeleton-based human action recognition. To exploit the implicit topology of multivariate synergy between joints, we introduce the Adaptive Non-uniform Hyper-graph, the Multi-head Hyper-graph Convolution and virtual connections. We carry out experiments on the dataset NTU-RGB+D 60 & 120, NW-UCLA to validate the effectiveness of Hyper-GCN. The experimental analysis verifies that our design can improve the recognition performance. To the best of our knowledge, Hyper-GCN achieves the SOTA performance on 3 public datasets. The involvement of adaptive non-uniform hyper-graph modeling indeed extends existing GCN-based action recognition paradigms.

### v3

@@ Abstract
The shared topology of human skeletons motivated the recent investigation of graph convolutional network (GCN) solutions for action recognition. However, most of the existing GCNs rely on the binary connection of two neighboring vertices (joints) formed by an edge (bone), overlooking the potential of constructing multi-vertex convolution structures. Although some studies have attempted to utilize hyper-graphs to represent the topology, they rely on a fixed construction strategy, which limits their adaptivity in uncovering the intricate latent relationships within the action. In this paper, we address this oversight and explore the merits of an adaptive hyper-graph convolutional network (Hyper-GCN) to achieve the aggregation of rich semantic information conveyed by skeleton vertices. In particular, our Hyper-GCN adaptively optimises the hyper-graphs during training, revealing the action-driven multi-vertex relations. Besides, virtual connections are often designed to support efficient feature aggregation, implicitly extending the spectrum of dependencies within the skeleton. By injecting virtual connections into hyper-graphs, the semantic clues of diverse action categories can be highlighted. The results of experiments conducted on the NTU-60, NTU-120, and NW-UCLA datasets demonstrate the merits of our Hyper-GCN, compared to the state-of-the-art methods. The code is available at https://github.com/6UOOON9/Hyper-GCN.
1 Introduction
Skeleton-based human action recognition is a popular research topic in artificial intelligence, with practical applications in video understanding, video surveillance, human-computer interaction, robot vision, VR and AR [28, 29, 16, 30, 49, 49, 42, 21]. In general, a skeleton sequence contains a series of 2D or 3D coordinates, which can easily be collected by low-cost depth sensors or obtained by video-based pose estimation algorithms [35, 26, 34, 43]. Compared to RGB and optical flow images, skeleton data, which represents the basic physical structure of a human being, is of lower dimension, conveying human action with higher efficiency. Moreover, it is robust to illumination changes and scene variations. For these reasons, the adoption of this structural data is very popular in skeleton-based action recognition [10, 45, 25, 35, 6, 3, 51].
@@ 1 Introduction
Figure 2: The framework of our Hyper-GCN. Part (a) represents the Multi-head Hyper-graph Convolution (M-HGC) module. Part (b) represents the process of constructing an adaptive hyper-graph. Part (c) represents the architecture of Hyper-GCN. $\mathbf{F_{p}}$ represents the features of real joints. $\mathbf{F}_{h}$ represents the hyper-joints which are learnable parameters. $\hat{A}$ represents the physical topology.
2 Related Work
2.1 Graph Topology for Action Recognition
@@
2.2 Feature Configurations for Action Recognition
@@ 3.1 Preliminaries
3.2 Hyper-graph
Here, we use $\mathcal{G_{H}}=(\mathcal{V_{H}},\mathcal{E_{H}},\mathcal{W_{H}})$ to define the spatial hyper-graph [12] with human skeleton. $\mathcal{V_{H}}$ and $\mathcal{E_{H}}$ follow similar definitions in the normal graph, which represent the set of joints and the set of hyper-edges. In addition, we introduce $\mathcal{W_{H}}$ to represent the weights of each hyper-edge. Since a hyper-edge contains multiple joints, the corresponding topology can no longer be simply represented by an adjacency matrix. Therefore, we introduce the incidence matrix $H\in\mathbb{R}^{N\times M}$ to describe the topology of each joint in the hyper-graph. $N$ represents the number of joints and $M$ represents the number of hyper-edges. Given $v\in\mathcal{V_{H}}$ and $e\in\mathcal{E_{H}}$, the values of incidence matrix be determined by:
@@ 3.2 Hyper-graph
3.3 Adaptive Non-uniform Hyper-graph
@@
Given the general features $\mathbf{X}\in\mathbb{R}^{N\times C}$ of spatial skeleton joints, $N$ represent the number of joints, $C$ represent the channels. Our A-NHG construct the hyper-edge with each joint as the center of mass, and a total of $N$ hyper-edges are constructed. A-NHG utilises the Euclidean distance to measure the joint difference.
@@
$h_{i,j}$ =$=\begin{cases}\frac{\text{exp}(-m_{i,j})}{\sum_{k\in set_{i}}{\text{exp}(-m_{i,k})}},&j\in set_{i},\\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 0,&j\notin set_{i}.\end{cases}$ (8)
@@
3.4 Multi-head Hyper-graph Convolution
@@
We use separate branches to independently perform hyper-graph convolution on the topologies represented by the multi-head hyper-graphs, thereby enhancing the computational efficiency of the Hyper-GCN. For instance, the features $\mathbf{F_{in}}\in\mathbb{R}^{C_{\text{in}}\times T\times V}$ are divided into 888 separate branches along the channel dimension, which are processed by 888 M-HGC parallelly, delivering 888 separate hyper-graphs.
Next, we perform temporal pooling, and then the obtained spatial information is the evidence to determine the optimal hyper-graph construction. This can efficiently decouple the temporal and spatial clues. After performing temporal average-pooling, we can obtain $\mathbf{\bar{F}_{in}}\in\mathbb{R}^{C_{\text{in}}\times V}$.
The hyper-graph is depicted by the incidence matrix and weight matrix. So, we use a separate mapping function for each head to embed the features into the subspace for constructing the hyper-graph. This design enables preserving the original features to reflect their spatial relevance. After that, we introduce the A-NHG to obtain the incidence matrix. For the weight matrix, we adopt MLP to measure each hyper-edge. The hyper-graph obtained operation can be formulated as follows:
$H$ $=\uplus^{8}_{k=1}\text{A-NHG}(\mathbf{\bar{F}^{\mathnormal{k}}_{in}}),$ (9)
$W$ $=\sigma_{2}(\Psi_{2}(\uplus^{8}_{k=1}\sigma_{1}(\Psi^{k}_{1}(\mathbf{\bar{F}^{\mathnormal{k}}_{in}})))),$ (10)
@@
$\mathbf{F_{out}}$ $=\uplus^{8}_{k=1}{(\hat{A_{k}}+\alpha\cdot\hat{H_{k}})\mathbf{F^{\mathnormal{k}}_{in}}P_{k}},$ (11)
where $\hat{A}$ and $\hat{H}$ denote the normalized physical adjacency matrix and normalized incidence matrix. $P$ is the learnable weight parameter for feature transform. $\alpha$ is a learnable parameter for the topology fusion of each head.
3.5 Virtual Connections
@@
The shape of hyper-joints is consistent with the features of physical joints $\mathbf{F_{p}}\in\mathbb{R}^{C\times T\times V}$, which is described as $\mathbf{F_{h}}\in\mathbb{R}^{C\times T\times V_{h}}$ ($V_{h}$ represents the number of hyper-joints). To ensure alignment in the temporal dimension, the hyper-joints are shared across each frame. Specifically, the learning of hyper-joints is supervised by the loss function, which guides these hyper-joints to support generalisable driven features embedded within large amounts of data. Furthermore, we set independent hyper-joints at each layer of Hyper-GCN, aiming to harmonise features at different depths. Typically, these hyper-joints are involved in spatial hyper-graph convolution rather than in temporal convolution.
@@
$C=\frac{\mathbf{F_{h}F^{\mathnormal{T}}_{h}}}{||\mathbf{F_{h}}||^{2}},$ (12)
Specifically, in Divergence Loss, we measure the differences of hyper-joints with the mean of the cosine matrix $C$ in each layer. Since the correlation of hyper-joints themselves cannot be optimised, we manually eliminate this part by minus the $V_{h}$. The loss calculation can be formulated as:
@@
where $V_{h}$ represents the number of hyper-joints we introduced. $C_{l}$ represents the cosine matrix of the $l$-th layer. $\mathcal{L}_{\text{CE}}$ represents the cross-entropy loss. Additionally, we manually connect the hyper joints to all the physical joints, which optimises the topology from the datasets.
3.6 Entire Architecture
Hyper-GCN consists of an embedding layer [6], 9 spatial-temporal convolution layers as shown in Figure 2 (c). Each layer consists of the proposed M-HGC and Multi Scale Temporal convolution (MS-TC) [3]. 9 Layers are categorised into 3 stages. In each stage, we introduce the dense connections to integrate deep and shallow features in each stage. Additionally, to further validate the potential of Hyper-GCN, we propose a base version and a large version. The channels in each stage are set to 128, 256, 256 for the base version and 128, 256, 512 for the large version.
Category Methods Modalities Params (M) GFLOPs NTU-RGB+D 60 NTU-RGB+D 120 NW-UCLA (%)
X-Sub (%) X-View (%) X-Sub (%) X-Set (%)
GCN ST-GCN [52] J+B - - 81.5 88.3 70.7 73.2 -
2s-AGCN [36] J+B - - 88.5 95.1 82.5 84.2 -
DC-GCN+ADG [4] J+B+JM+BM 4.9 1.83 90.8 96.6 86.5 88.1 95.3
MS-G3D [27] J+B+JM+BM 2.8 5.22 91.5 96.2 86.9 88.4 -
MST-GCN [11] J+B+JM+BM 12.0 - 91.5 96.6 87.5 88.8 -
CTR-GCN [3] J+B+JM+BM 1.5 1.97 92.4 96.4 88.9 90.6 96.5
EfficientGCN-B4 [41] J+B+JM+BM 2.0 15.20 91.7 95.7 88.3 89.1 -
InfoGCN [6] J+B+JM+BM 1.6 1.84 92.7 96.9 89.4 90.7 96.6
FR Head [56] J+B+JM+BM 2.0 - 92.8 96.8 89.5 90.9 96.8
HD-GCN* [23] J+B+J’+B’ 1.7 1.77 93.0 97.0 89.8 91.2 96.9
DS-GCN [51] J+B+JM+BM - - 93.1 97.5 89.2 90.3 -
BlockGCN [57] J+B+JM+BM 1.3 1.63 93.1 97.0 90.3 91.5 96.9
Transformer DSTA-Net [38] J+B+JM+BM 3.5 16.18 89.5 95.7 86.6 89.0 -
IIP-Transformer [47] J+B+JM+BM 2.9 7.20 89.5 95.7 89.9 90.9 -
SkateFormer [8] J+B+JM+BM 2.0 3.62 93.5 97.8 89.8 91.4 98.3
HGCN Hyper-GNN [13] J+B+JM+BM - - 89.5 95.7 - - -
Selective-HCN [58] J+B+JM+BM - - 90.8 96.6 - - -
DST-HCN [48] J+B+JM+BM 3.5 2.93 92.3 96.8 88.8 90.7 96.6
Ours (B) J+B+JM+BM 1.1 1.63 93.3 97.4 90.5 91.7 97.2
Ours (L) J+B+JM+BM 2.3 2.88 93.7 97.8 90.9 92.0 97.6
Table 1: Comparison of Hyper-GCN with advanced solutions on NTU-RGB+D 60, NTU-RGB+D 120, and NW-UCLA datasets. For a fair comparison, we use a 4-streams ensemble to evaluate the performance of the mentioned methods. J, B, JB, and JM represent the joint, bone, joint motion, and bone motion. In addition, for HD-GCN [23], we use the results from 4-streams according to the original papers for a fair comparison. The bold font represents the best result and underline font represents the 2-nd best result.
4 Evaluation
4.1 Datasets
@@
4.2 Implementation Details
Our training and evaluation stages are on a single GPU RTX 3090. In the training phase, Hyper-GCN is optimised by Stochastic Gradient Descent (SGD) with Nesterov momentum set at 0.9 and a weight decay at 0.0004. Our implementation uses label smooth cross-entropy loss with the Divergence Loss we proposed. We set a total of 140 epochs with the start 5 warm-up epochs. The initial learning rate is 0.05, which is reduced to 0.005 at epoch 110 and to 0.0005 at epoch 120.
4.3 Comparison with the State-of-the-Art
Multi-stream ensemble proposed in  [39] have been proven effective by most of the existing state-of-the-art methods. We also use a 4-stream ensemble to evaluate the base version and the large version of Hyper-GCN. The detailed results are reported in Table 1.
In summary, the base version of Hyper-GCN comprehensively outperforms all GCN-based and HGCN-based (hyper-graph utilised) SOTA and surpasses the Transformer-based SOTA on the NTU120, with a most lightweight design. Moreover, when scaled to match the parameters of the lightest Transformer-based method, the large version of Hyper-GCN achieves 1-st place in 4 benchmarks and 2-nd place in 1 benchmark. This demonstrates the effectiveness and merits of Hyper-GCN.
4.4 Ablation Study
We conduct ablation experiments with visualisations to analyse the effectiveness of our design. redAll the ablation study is evaluated with the joint modality of base version.
@@ 4.4 Ablation Study
Table 2: Ablation of the hyper-parameter $K$ in A-NHG. The Uniform represents K-uniform hyper-graph, in which one hyper-edge contains K joints. The non-uniform represents non-uniform hyper-graph, in which one joint is contained by $K$ hyper-edge.
Figure 4: Visualisation of hyper-graph in selected actions. For ease of observation, the joints contained in the selected hyper-edge are highlighted in red.
@@
The results are reported in Table 2. In both uniform and non-uniform hyper-graphs, a very large $K$ means that more joints are contained in each hyper-edge. Even though these ”extra” joints may be optimised for minimal weight, they are likely to act as ”noise”, disrupting the information interaction represented by the hyper-edge. As the $K$ decreases, it is difficult for the hyper-edge to represent complex joint combination patterns. Therefore, setting $K=5$ in uniform and $K=9$ in Non-uniform achieves the best performance. The non-uniform hyper-graph of A-NHG outperforms the uniform hyper-graph, as the combination of skeleton joints in different actions is diverse.
Furthermore, we visualise the hyper-graph inferred by Hyper-GCN on selected actions in Figure 4. The hyper-graph construction focuses on the joints that are most relevant to the corresponding action category. For “Kicking something” as part (d), the left leg joints and the right hand joints are modelled as a hyper-edge. Because in this action, the person’s right hand is previously extended forward, and the left leg swings relatively to the right hand. Similarly, in the action ”Stand up” as shown in part (b), the hand and foot joints serve as fixed joints, while the remaining body joints move relative to them. In addition, ”Taking a selfie” and ”Drinking water” both focus on the main joints of movement with the above view as Figure 4 (a), (c).
@@
w/o M-HGC w/ 1 84.9 (0.2$\uparrow$ 0.2↑ 0.2) 84.9 (0.2$\uparrow$ 0.2↑ 0.2)
w/ 3 84.9 (0.2$\uparrow$ 0.2↑ 0.2) 85.2 (0.5$\uparrow$ 0.5↑ 0.5)
w/ 5 84.7 85.0 (0.3$\uparrow$ 0.3↑ 0.3)
w M-HGC w/ 1 86.7 (2.0$\uparrow$ 2.0↑ 2.0) 86.7 (2.0$\uparrow$ 2.0↑ 2.0)
w/ 3 86.6 (1.8$\uparrow$ 1.8↑ 1.8) 86.9 (2.2$\uparrow$ 2.2↑ 2.2)
w/ 5 86.6 (1.8$\uparrow$ 1.8↑ 1.8) 86.8 (2.1$\uparrow$ 2.1↑ 2.1)
Table 3: Ablations on hyper joints. w/ $N$ represents the number of hyper joints in each layer. w/ $\mathcal{L}_{h}$ represents the training with Divergence Loss. The hyper-parameter $K$ of M-HGC is set to 999.
@@
Virtual Connections: We conduct the ablation study to analyse the hyper joints and the Divergence Loss. The results are listed in Table 3. By comparison, the performance is not monotonically increasing with the number of hyper joints. The case of introducing only 333 hyper joints achieves the best performance. As our hyper joints are learned from a large amount of data, involving a large number of hyper joints can introduce redundant and ambiguous clues, degrading the performance.
@@
5 Conclusion
In this paper, we propose an Adaptive Hyper-graph Convolutional Network for skeleton-based human action recognition. To exploit the implicit topology of multivariate synergy between joints, we introduce the Adaptive Non-uniform Hyper-graph, the Multi-head Hyper-graph Convolution and virtual connections. We carry out experiments on the dataset NTU-RGB+D 60 & 120, NW-UCLA to validate the effectiveness of Hyper-GCN. The experimental analysis verifies that our design can improve the recognition performance. To the best of our knowledge, Hyper-GCN achieves the SOTA performance on 3 public datasets. The involvement of adaptive non-uniform hyper-graph modelling indeed extends existing GCN-based action recognition paradigms.
Acknowledgment This work was supported in part by the National Natural Science Foundation of China (62020106012, 62106089, 62332008, 62336004), the Engineering and Physical Sciences Research Council (EPSRC), U.K. (Grants EP/V002856/1, and EP/T022205/1), and the Fundamental Research Funds for the Central Universities (JUSRP202504007).